### v2

@@ 5 Properties of the Gale demand correspondence
    for any $\mathbf{q}\in\operatorname{pdom}(u)$, $\max_{y\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}b\log u(y)-\left\langle\mathbf{q},y\right\rangle<+\infty$;
@@
    $\{\mathbf{q}\leavevmode\nobreak\ |\leavevmode\nobreak\ \mathbf{q}>0\}\subseteq\operatorname{pdom}(u)$;
@@
    for any $\mathbf{p}\geq 0$, if $\max_{\left\langle\mathbf{p},y\right\rangle\leq b}\log u(y)<+\infty$, then $\mathbf{p}\in\operatorname{pdom}(u)$.
@@
    for any $M>0$, if there is a convergent sequence of prices $\mathbf{p}^{(k)}\in\operatorname{pdom}(u)$ with $\max_{y\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}b\log u(y)-\left\langle\mathbf{p},y\right\rangle<M$, $k=1,2,\ldots$ with a limit point $\mathbf{p}^{(k)}\to\bar{\mathbf{p}}$, then $\bar{\mathbf{p}}\in\operatorname{pdom}(u)$.
@@
We first prove property (i). According to the definition of $\operatorname{pdom}(u)$, there exists $b^{\prime}>0$ such that $\max_{y\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}b^{\prime}\log u(y)-\left\langle\mathbf{q},y^{\prime}\right\rangle<+\infty$. For any $b>0$ and for any $y\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$, if $\left\langle\mathbf{q},y\right\rangle\leq b$, then $b\log u(y)-\left\langle\mathbf{q},y\right\rangle<\frac{b}{b^{\prime}}\left(b^{\prime}\log u(y)-\left\langle\mathbf{q},y\right\rangle\right)+\frac{b-b^{\prime}}{b^{\prime}}\ \left\langle\mathbf{q},y\right\rangle$, which is bounded. Then, the result follows using Lemma 5.1.
Property (ii) holds as, by Lemma 5.1, $\max_{y\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}b\log u(y)-\left\langle\mathbf{q},y\right\rangle\leq\max_{\left\langle\mathbf{q},x\right\rangle\leq b}b\log u(x)-\left\langle\mathbf{q},x\right\rangle$, which is bounded as $x$ is bounded: $x_{j}\leq\frac{b}{q_{j}}<\infty$.
To prove property (iii), we let $\alpha=\min_{q_{j}>0}\{q^{\prime}_{j}/q_{j}\}$. Then, $\max_{y\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}(\alpha b_{i})\log u_{i}(y_{i})-\left\langle\mathbf{q}^{\prime},y_{i}\right\rangle\leq\max_{y\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}(\alpha b_{i})\log u_{i}(y_{i})-\alpha\left\langle\mathbf{q},y_{i}\right\rangle<+\infty$.
Property (iv) holds as, for any $y$ such that $\left\langle\mathbf{p},y\right\rangle\leq b$, $b\log u(y)-\left\langle\mathbf{q},y\right\rangle\leq b\log u(y)<+\infty$.
Finally, let us prove (v). For a contradiction, assume $\bar{\mathbf{p}}\notin\operatorname{pdom}(u)$, that is, $\sup_{y\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}b\log u(y)-\left\langle\bar{\mathbf{p}},y\right\rangle=\infty$. Consider a vector $y\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ with $b\log u(y)-\left\langle\bar{\mathbf{p}},y\right\rangle>2M$. Clearly, $y\neq 0$. Let us select $k$ such that $\|\mathbf{p}^{(k)}-\bar{\mathbf{p}}\|<M/\|y\|$. Then, $b\log u(y)-\left\langle\mathbf{p}^{(k)},y\right\rangle=b\log u(y)-\left\langle\bar{\mathbf{p}},y\right\rangle+\left\langle\bar{\mathbf{p}}-\mathbf{p}^{(k)},y\right\rangle>M$, in contradiction with the assumption on the sequence $\mathbf{p}^{(k)}$. ∎
We next prove Lemma 2.4. See 2.4

### v3

@@ 5 Agent-wise utility guarantee
Let $u(\cdot)$ be a $\Sigma$-Gale-substitutes utility and let $\mathbf{x}\in{\mathcal{D}}^{u}(\mathbf{p},b)$ for $\mathbf{p}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ and $b>0$. Then, $\mathbf{p}\in\operatorname{pdom}(u)$, and there exists $\mathbf{y}\in\mathcal{GD}^{u}(\mathbf{p},b)$ such that $\mathbf{y}\leq\mathbf{x}$.
We now restate Theorem 1.3, and sketch the proof. The full proof is given in Appendix B. See 1.3
Proof sketch.
We start with a competitive equilibrium $((\mathbf{x}_{i})_{i\in{\mathcal{A}}},\mathbf{p})$, and show that the dual of the Eisenberg–Gale program (5), $\varphi(\mathbf{q})$, has an optimal solution $\mathbf{q}^{\star}$ such that $\mathbf{q}^{\star}\leq\mathbf{p}$. The statement then follows from Lemma 5.1.
An intuitive argument is as follows: starting from $\mathbf{q}=\mathbf{p}$, where the total Gale demand is less than the supply (by Lemma 5.2), we decrease the prices in a tâtonnement-like procedure Cheung et al. (2020) using the Gale-substitutes property. As long as $\mathbf{q}$ is not a minimizer of $\varphi(.)$, there will be an underdemanded good $j$ with respect to Gale demands. Reducing $q_{j}$ for such a good $j$ may only decrease the Gale demand for all agents on all other goods. This iterative process converges to an equilibrium point and, since the prices are only decreasing, we can conclude that the optimal limit price vector satisfies $\mathbf{q}^{\star}\leq\mathbf{p}$.
Rather than analysing this continuous procedure, the proof in Appendix B establishes $\mathbf{q}^{\star}\leq\mathbf{p}$ directly, using Lagrangian duality. For any price vector $\mathbf{q}$, we get a subgradient in $\partial^{\vee}\varphi(\mathbf{q})$ from the difference between the supply and the total Gale demand. Further, for $\mathbf{q}\leq\mathbf{p}$ we get that if $q_{j}=p_{j}$ then the $j$-th component of this subgradient at $\mathbf{q}$ is non-negative. This is shown in Lemma B.1, by making use of an allocation $\mathbf{y}$ derived using Lemma 5.2 for the prices $\mathbf{q}$. Thus, in the region $\{\mathbf{q}\penalty 10000\ |\penalty 10000\ \mathbf{0}\leq\mathbf{q}\leq\mathbf{p}\}$, the subgradients on the boundary all point outward. Consequently, if $\mathbf{q}^{\star}\not\leq\mathbf{p}$, then we can take the projection $\mathbf{q}$ of $\mathbf{q}^{\star}$ to the boundary of the region, where this subgradient shows that $\mathbf{q}$ is also optimal. ∎
The following example shows the bound in Theorem 1.3 is tight.
Example 5.3.
Consider a market with $n$ agents and $n$ goods, where all agents have bounded linear utilities. Agent 1’s utility is
$u_{1}(\mathbf{x}_{1})=\sum_{j}\min\{1,x_{1j}\},$
while, for all other agents ($2\leq i\leq n$), they all have identical preferences on all goods except for the first one
$u_{i}(\mathbf{x}_{i})=\sum_{j>1}\min\{1,x_{ij}\}.$
In the Nash welfare maximizing allocation, each agent receives a bundle yielding utility exactly equal to $1$. However, consider a competitive equilibrium in which the price of the first good is 0, and the prices of all other goods are set to $\frac{n}{n-1}$. In this equilibrium, agent $1$ obtains the entire first good and a $\frac{n-1}{n}$ fraction of the other goods, leading to a total utility of
$u_{1}=1+\frac{n-1}{n},$
which approaches $2$ as $n\to\infty$.
Remark 5.4.
The assumption that the utilities are $\Sigma$-Gale-substitutes is important in Theorem 1.3. Example E.5 demonstrates that for non-Gale substitute utilities, some agents may achieve significantly higher utility in a competitive equilibrium than in the Nash welfare maximizing allocation.
Remark 5.5.
Theorem 1.3 naturally raises a question about the converse: for any agent $i$, does there always exist a CE $(\mathbf{x}_{i})_{i\in{\mathcal{A}}}$ such that $u_{i}(\mathbf{x}_{i})\geq C\cdot u_{i}(\mathbf{y}_{i})$ for some constant $C>0$, where $(\mathbf{y}_{i})_{i\in{\mathcal{A}}}$ is a Nash welfare maximizing allocation? We show that such a constant-factor guarantee is not always possible by presenting an example (see Example E.4) in which the CE is unique, yet there exists an agent whose utility under this equilibrium is significantly lower than in the Nash welfare maximizing allocation.
6 Gale substitute utilities
First, we show that separable utilities are $\Sigma$-Gale-substitutes. See 1.4 For this, we first adapt Lemma 4.2 to the setting of separable utilities.
Lemma 6.1.
Let $u(\mathbf{x})=\sum_{j\in{\mathcal{M}}}v_{j}(x_{j})$ be a separable utility function. For any $b>0$, $\mathbf{q}\in\operatorname{pdom}(u)$, $\mathbf{y}\in\mathcal{GD}^{u}(\mathbf{q},b)$, we have $u(\mathbf{y})>0$. Further, $\mathbf{y}\in\mathcal{GD}^{u}(\mathbf{q},b)$ if and only if there exist supergradients $g_{j}\in\partial^{\wedge}v_{j}(y_{j})$, $j\in{\mathcal{M}}$ satisfying
  1. (i)
    $g_{j}\geq 0$; and
  2. (ii)
    $g_{j}\leq q_{j}u(\mathbf{y})/b$ with equality whenever $y_{j}>0$.
We now show that strictly concave and differentiable utility functions satisfy Definition 2.14 (i) : for any $b>0$, any $\mathbf{q},\mathbf{q}^{\prime}\in\operatorname{pdom}(u)$, and any $\mathbf{y}\in\mathcal{GD}^{u}(\mathbf{q},b)$, if $\mathbf{q}\geq\mathbf{q}^{\prime}$, then there exists a $\mathbf{y}^{\prime}\in\mathcal{GD}^{u}(\mathbf{q}^{\prime},b)$ such that $y^{\prime}_{j}\leq y_{j}$ whenever $q_{j}^{\prime}=q_{j}$.
Proof of Theorem 1.4 for strictly concave and differentiable utilities.
Let $u(\mathbf{y})=\sum_{j\in{\mathcal{M}}}v_{j}(y_{j})$ be a separable and differentiable concave utility function, and let $\nabla_{\!j}u(\mathbf{y})=\partial v_{j}(y_{j})/\partial y_{j}$ denote the $j$-th component of the gradient. Recall that for differentiable utilities, the only supergradient of $v_{j}(y_{j})$ is $\nabla_{\!j}u(\mathbf{y})$. By Lemma 6.1, $\nabla_{\!j}u(\mathbf{y})\leq q_{j}u(\mathbf{y})/b$, with equality whenever $y_{j}>0$. Let $\mathbf{y}\in\mathcal{GD}^{u}(\mathbf{q},b)$ and $\mathbf{y}^{\prime}\in\mathcal{GD}^{u}(\mathbf{q}^{\prime},b)$ for $\mathbf{q}^{\prime}\leq\mathbf{q}$.
We first claim that $u(\mathbf{y}^{\prime})\geq u(\mathbf{y})$. For a contradiction, assume $u(\mathbf{y}^{\prime})<u(\mathbf{y})$. From the above, whenever $y_{j}>0$, we have
$\nabla_{\!j}u(\mathbf{y})=q_{j}u(\mathbf{y})/b\geq q^{\prime}_{j}u(\mathbf{y}^{\prime})/b\geq\nabla_{\!j}u(\mathbf{y}^{\prime})\,.$
By strict concavity, this means $y^{\prime}_{j}\geq y_{j}$. Thus, we see that $\mathbf{y}^{\prime}\geq\mathbf{y}$, a contradiction to the monotonicity of $u(.)$ and the assumption $u(\mathbf{y}^{\prime})<u(\mathbf{y})$.
From $u(\mathbf{y}^{\prime})\geq u(\mathbf{y})$, for every good $j$ with $q_{j}=q^{\prime}_{j}$ and $y_{j}>0$, $\nabla_{\!j}u(\mathbf{y}^{\prime})\geq\nabla_{\!j}u(\mathbf{y})$ follows. By strict concavity, we must have $y^{\prime}_{j}\leq y_{j}$ as required. ∎
The proof without the differentiability and strict concavity assumptions poses additional difficulties: for example, we would not be able to conclude $y^{\prime}_{j}\leq y_{j}$ in the last step, and might need a specific choice of $y^{\prime}_{j}$ from $\mathcal{GD}^{u}(\mathbf{q}^{\prime},b)$. The proof of the general case is given in Appendix C.
Our final theorem is the Gale substitutability of generalized network utility functions. This is deferred to Appendix D, and we only present a high level summary here. See 1.5
To verify the Gale substitutes property, we need to take two price vectors $\mathbf{q}\geq\mathbf{q}^{\prime}$ and $\mathbf{y}\in\mathcal{GD}^{u}(\mathbf{q},b)$, and construct a $\mathbf{y}^{\prime}\in\mathcal{GD}^{u}(\mathbf{q^{\prime}},b)$ such that $y^{\prime}_{j}\leq y_{j}$ whenever $q^{\prime}_{j}=q_{j}$; we have a second, analogous construction for budget changes. The proof is a finite algorithm that transforms the optimal generalized flow for demands $\mathbf{y}$ to $\mathbf{y}^{\prime}$ such that we throughout maintain the primal-dual slackness condition corresponding to the Gale demand correspondence. The algorithm builds on the literature on maximum generalized flow algorithms. In particular, we use ideas from the classical highest gain augmenting path algorithm by Onaga Onaga (1966). While this does not give a polynomial-time algorithm, it was the starting point of further such algorithms, including Goldberg et al. (1991), Olver and Végh (2020), Végh (2014).
Although generalized network utilities are not $\Sigma$-Gale-substitutes in general (see Example E.6), we can still establish the following theorem with a slight modification to the proof of Theorem 1.3; the proof is also given in Appendix D.
See 1.6
7 Discussion
In this paper, we compared competitive equilibrium and Nash welfare maximizing solutions in the Fisher market setting. We identified Gale-substitutes as a broad family of utility functions where the Nash welfare uniformly guarantees every agent at least half of their maximum utility at any competitive equilibrium. We also introduced generalized network utilities. This is a very general construction with rich combinatorial structure and may be a relevant example in various market and fair division settings.
Gale-substitutes may form a compelling subclass of concave functions worth investigating on its own. In the context of indivisible goods, the analogous class of gross substitute utilities, introduced by Kelso and Crawford (1982), turned out to be the same as the important class of $M^{\natural}$-concave utilities in discrete convex analysis Fujishige and Yang (2003). In particular, one may wonder whether the Gale-substitutes property implies submodularity, that is, $u(\mathbf{x})+u(\mathbf{y})\geq u(\mathbf{x}\vee\mathbf{y})+u(\mathbf{x}\wedge\mathbf{y})$ for any $\mathbf{x},\mathbf{y}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$, where $\mathbf{x}\vee\mathbf{y}$ and $\mathbf{x}\wedge\mathbf{y}$ denote the pointwise maximum and minimum of the two vectors, respectively. This property trivially holds for separable utilities and can also be shown for generalized network utilities. Another interesting question is whether computing competitive equilibrium in generalized network utilities lies in PPAD.
We note that one can construct an example of a concave quadratic function that is submodular but not Gale-substitutes. On the other hand, it is unclear whether Theorem 1.6 would remain true for submodular utilities instead of Gale-substitutes. One would also be interested in whether Theorem 1.6 holds for matching utilities, which, like submodular utilities, are not necessarily Gale-substitutes (see Example E.7). Another natural question related to the well-studied notion of weak gross substitutes (WGS) utilities: Gale-substitutes includes utility functions that are not WGS, such as concave separable; but it is not clear whether all WGS utilities are also Gale-substitutes.
Gale-substitutes utilities may also have potential algorithmic applications. The WGS property leads to natural tâtonnement and auction mechanisms for finding competitive equilibria, see, e.g., Arrow et al. (1959), Arrow and Hurwicz (1960), Bei et al. (2019a), Codenotti et al. (2005b, c), Garg et al. (2023a). In particular, Garg et al. (2023a) gives an auction algorithm for maximizing Nash welfare under capped SPLC utilities. Similar methods could be extended to the broader class of Gale substitutes utilities. The Gale-substitutes property may be useful to design simple such mechanisms and combinatorial algorithms for solving the convex program (EG). For the linear case, Devanur et al. (2008) gave a primal-dual algorithm that was followed by a series of works on combinatorial algorithms for various market settings; see, e.g., Chaudhury et al. (2022), Cole and Gkatzelis (2015), Duan and Mehlhorn (2015), Garg and Végh (2023).
Appendix A Properties of the Gale demand correspondence
In this Appendix, we prove Lemma 2.4. We also derive some technical properties of the price domain $\operatorname{pdom}(u)$ that will be used in the subsequent Appendices.
See 2.4
Proof.
Let $M\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\sup_{\mathbf{y}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}b\log u(\mathbf{y})-\left\langle\mathbf{p},\mathbf{y}\right\rangle$, which implies the existence of a sequence of allocations $\{\mathbf{y}^{(k)}\}_{k=1}^{\infty}$ such that $b\log u(\mathbf{y}^{(k)})-\left\langle\mathbf{p},\mathbf{y}^{(k)}\right\rangle$ converges to $M$. We note that the finiteness of $M$ follows from Lemma 4.1, which is explicitly established in Lemma A.1(i). We show the existence of an allocation $\mathbf{y}^{*}$ satisfying $b\log u(\mathbf{y}^{*})-\left\langle\mathbf{p},\mathbf{y}^{*}\right\rangle=M$.
According to Lemma 4.1, there exists a sequence of $\bar{\mathbf{y}}^{(k)}$ such that $b\log u(\bar{\mathbf{y}}^{(k)})-\left\langle\mathbf{p},\bar{\mathbf{y}}^{(k)}\right\rangle$ converges to $M$, with $0\leq\left\langle\mathbf{p},\bar{\mathbf{y}}^{(k)}\right\rangle\leq b$ for all $k$. Thus, we can select a subsequence $\bar{\mathbf{y}}^{(i_{k})}$ such that $b\log u(\bar{\mathbf{y}}^{(i_{k})})-\left\langle\mathbf{p},\bar{\mathbf{y}}^{(i_{k})}\right\rangle$ converges to $M$ and $\left\langle\mathbf{p},\bar{\mathbf{y}}^{(i_{k})}\right\rangle$ converges to $b^{\prime}$ for some $b^{\prime}\in[0,b]$. Consequently, $b\log u(\bar{\mathbf{y}}^{(i_{k})})$ converges to $M+b^{\prime}$ as $\left\langle\mathbf{p},\bar{\mathbf{y}}^{(i_{k})}\right\rangle$ converges to $b^{\prime}$. By Assumption 1, there exists an allocation $\mathbf{y}^{*}$ such that $b\log u(\mathbf{y}^{*})=M+b^{\prime}$ and $\left\langle\mathbf{p},\mathbf{y}^{*}\right\rangle\leq b^{\prime}$. This completes the proof. ∎
Lemma A.1.
@@ Appendix A Properties of the Gale demand correspondence
    for any $\mathbf{q}\in\operatorname{pdom}(u)$, $\max_{\mathbf{y}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}b\log u(\mathbf{y})-\left\langle\mathbf{q},\mathbf{y}\right\rangle<+\infty$;
@@
    $\{\mathbf{q}\penalty 10000\ |\penalty 10000\ \mathbf{q}>0\}\subseteq\operatorname{pdom}(u)$;
@@
    for any $\mathbf{p}\geq 0$, if $\max_{\left\langle\mathbf{p},\mathbf{y}\right\rangle\leq b}\log u(\mathbf{y})<+\infty$, then $\mathbf{p}\in\operatorname{pdom}(u)$.
@@
    for any $M>0$, if there is a convergent sequence of prices $\mathbf{p}^{(k)}\in\operatorname{pdom}(u)$ with $M>\max_{\mathbf{y}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}b\log u(\mathbf{y})-\left\langle\mathbf{p}^{(k)},\mathbf{y}\right\rangle$, $k=1,2,\ldots$ with a limit point $\mathbf{p}^{(k)}\to\bar{\mathbf{p}}$, then $\bar{\mathbf{p}}\in\operatorname{pdom}(u)$.
@@
We first prove property (i). According to the definition of $\operatorname{pdom}(u)$, there exists $b^{\prime}>0$ such that $\max_{\mathbf{y}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}b^{\prime}\log u(\mathbf{y})-\left\langle\mathbf{q},\mathbf{y}\right\rangle<+\infty$. For any $b>0$ and for any $\mathbf{y}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$, if $\left\langle\mathbf{q},\mathbf{y}\right\rangle\leq b$, then $b\log u(\mathbf{y})-\left\langle\mathbf{q},\mathbf{y}\right\rangle<\frac{b}{b^{\prime}}\left(b^{\prime}\log u(\mathbf{y})-\left\langle\mathbf{q},\mathbf{y}\right\rangle\right)+\frac{b-b^{\prime}}{b^{\prime}}\ \left\langle\mathbf{q},\mathbf{y}\right\rangle$, which is bounded. Then, the result follows using Lemma 4.1.
Property (ii) holds as, by Lemma 4.1, $\max_{\mathbf{y}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}b\log u(\mathbf{y})-\left\langle\mathbf{q},\mathbf{y}\right\rangle\leq\max_{\left\langle\mathbf{q},\mathbf{x}\right\rangle\leq b}b\log u(\mathbf{x})-\left\langle\mathbf{q},\mathbf{x}\right\rangle$, which is bounded as $\mathbf{x}$ is bounded: $x_{j}\leq\frac{b}{q_{j}}<\infty$.
To prove property (iii), we let $\alpha=\min_{q_{j}>0}\{q^{\prime}_{j}/q_{j}\}$. Then, $\max_{\mathbf{y}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}(\alpha b)\log u(\mathbf{y})-\left\langle\mathbf{q}^{\prime},\mathbf{y}\right\rangle\leq\max_{\mathbf{y}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}(\alpha b)\log u(\mathbf{y})-\alpha\left\langle\mathbf{q},\mathbf{y}\right\rangle<+\infty$.
Property (iv) holds as, for any $\mathbf{y}$ such that $\left\langle\mathbf{p},\mathbf{y}\right\rangle\leq b$, $b\log u(\mathbf{y})-\left\langle\mathbf{p},\mathbf{y}\right\rangle\leq b\log u(\mathbf{y})<+\infty$.
Finally, let us prove (v). For a contradiction, assume $\bar{\mathbf{p}}\notin\operatorname{pdom}(u)$, that is, $\sup_{\mathbf{y}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}b\log u(\mathbf{y})-\left\langle\bar{\mathbf{p}},\mathbf{y}\right\rangle=\infty$. Consider a vector $\mathbf{y}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ with $b\log u(\mathbf{y})-\left\langle\bar{\mathbf{p}},\mathbf{y}\right\rangle>2M$. Clearly, $\mathbf{y}\neq 0$. Let us select $k$ such that $\|\mathbf{p}^{(k)}-\bar{\mathbf{p}}\|<M/\|\mathbf{y}\|$. Then, $b\log u(\mathbf{y})-\left\langle\mathbf{p}^{(k)},\mathbf{y}\right\rangle=b\log u(\mathbf{y})-\left\langle\bar{\mathbf{p}},\mathbf{y}\right\rangle+\left\langle\bar{\mathbf{p}}-\mathbf{p}^{(k)},\mathbf{y}\right\rangle>M$, in contradiction with the assumption on the sequence $\mathbf{p}^{(k)}$. ∎
Appendix B Efficiency of Nash welfare for Gale-substitutes
In this Appendix, we give the full proof of Theorem 1.3. We show that for $\Sigma$-Gale substitutes utility functions, any competitive equilibrium $((\mathbf{x}_{i})_{i\in{\mathcal{A}}},\mathbf{p})$ admits Gale prices $\mathbf{q}^{\star}$ such that $\mathbf{q}^{\star}\leq\mathbf{p}$. The theorem statement then follows from Lemma 5.1. We start by re-stating and proving Lemma 5.2.
See 5.2